Displacement

You are driving home on a weekend from school 55 mi/h for 110 miles. it then starts to snow and you slow to 35 mi/h. you arrive home after driving 4 hours and 15minutes. How far is your hometown from school?
Answer #1
If you drive at 55 mi/h for 110 miles, this means you drovefor 110 / 55 = 2 hoursYour total driving time is 4 hours and 15 minutes which meansthat you drove 2 and 15 minutes at the rate of 35 mi/h2 hours and 15 minutes = 2*60 + 15 = 135minutes35 miles / 60 minutes * ( 135 minutes) = 78.75milesSo total miles driven is 78.75 + 110 =188.75 miles
answered by: raven braswell
Answer #2
55mi * 110mi =2h
h

35mi * 4.15h = 145.24
h

145.24 = 72.625mi
2
answered by: vicky
Answer #3
There is more than one way to approach this problem.First make sure all of the units are compatable. Youmust change 4 hr 15 min to hours. Since 15 min is 1/4 of anhour you can write the timeas 4.25 hr.In the first part of the situation you are driving 55 mi/h for110 mi, using the average speed equation:v=d/t you can solve for the time required: t = d/vt = 110 mi/55 mi/h = 2 hoursThis means you drive 2 hours before the snow and 4.25 hourstotal, so 2.25 hours are driven in the snow at 35 mi/h.Go back to the average speed equation: v=d/t and solvefor displacement:d = vtd = (35 mi/h) (4.25 h) = 78.75 miYou had already driven 110 miles, so the total distanceis110 mi + 79 mi = 189 miBe careful: This question is asking aboutdistance,not displacement!!
answered by: rehmanullah